OpenStudy (anonymous):

What is the y-intercept in the equation y = x – 5?

ganeshie8 (ganeshie8):

y = mx + c y-intercept = c

ganeshie8 (ganeshie8):

y = x – 5 y-intercept = ?

OpenStudy (anonymous):

5?

ganeshie8 (ganeshie8):

dont neglect the signs :)

ganeshie8 (ganeshie8):

y = x - 5 y = x + (-5) c = -5 so y-intercept is -5

OpenStudy (anonymous):

negative 5 lol okay

ganeshie8 (ganeshie8):

yes :)